OpenStudy (anonymous):

A firecracker shoots up from a hill 140 feet high, with an initial speed of 100 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? Five seconds Seven seconds Nine seconds 11 seconds

OpenStudy (anonymous):

@ganeshie8 @Abhisar

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

yeah I already did that. h(t) = -16t^2 + 100t + 140

OpenStudy (dantefemboy):

D.

OpenStudy (anonymous):

I am pretty sure it is not seven seconds though

OpenStudy (anonymous):

and @DanteFemboy do you know why its D?

OpenStudy (perl):

then set it equal to zero

OpenStudy (perl):

solve h(t) =0 -16t^2 + 100t + 140= 0

OpenStudy (perl):

you have to use quadratic formula

OpenStudy (perl):

t = 7.42 , -1.18 we can reject the negative solution

OpenStudy (dantefemboy):

If my calculator is right i solved it to be 140. And what goes into 140 evenly? 11.83215957. So mostly 11. Is that right @perl if not then I need to update my stuff.

OpenStudy (dantefemboy):

I think i'm wrong. Big time lol.

OpenStudy (perl):

you must set it equal to zero :)

OpenStudy (dantefemboy):

But you have to move the X to one side right?

OpenStudy (anonymous):

One second \[t= −b (+ or −) (√b2−4ac)\div2a\] right?

OpenStudy (perl):

you mean t, it is already moved to one side

OpenStudy (dantefemboy):

Well yeah t. Ah I know what I messed up on lol

OpenStudy (anonymous):

ok so t = 7.4 t = -1.17

OpenStudy (anonymous):

h(t) = -16(7.4)^2 + 100(-1.17) + 140 right?

OpenStudy (perl):

and you shouldnt plug in different t values, they are separate solutions

OpenStudy (anonymous):

I'm confused, since I found t, what do I do with it now?

OpenStudy (perl):

youre done, now you know how long it will take to hit the ground

OpenStudy (perl):

you have your time. but the negative solution you can ignore

OpenStudy (anonymous):

wait I think I get it now. So, it will take 7 seconds?

OpenStudy (perl):

yes

OpenStudy (anonymous):

Oh! Sorry I am usually good with this but I can't think today and I am extremely tired. Lol thank you! :)

OpenStudy (perl):

:)